Find the value of 14 C 1 + 14 C3 + 14 C5 + ........ 14 C11

Grade:

1 Answers

Arun
5 years ago
 
14C1 + 14C3 + 14C5 + ….........+ 14C11 + 14C13 = 2^14 / 2 = 2^13
 
14C1 + 14C3 + 14C5 + …......... + 14C11 = 2^13 – 14C13 = 2^13 – 14
 
In case of any difficulty, please feel free to ask.
 
Regards
Arun (askIITians forum expert)